# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolic Fuzzing

### v2

@@ 5 Found Exploits
Running mpfuzz on six leading Ethereum execution-layer clients, namely, Geth, Besu, Nethermind, Erigon, Reth, and OpenEthereum, leads to the discovery of $22$ bugs on all six Ethereum clients, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions. We describe two selected patterns of these bugs as follows.
5.1 Found Eviction Attacks (Selected)
Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high fees to evict normal transactions initially stored in the mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions. 2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high fee, of high value $v_{1}$, and from the same $k$ accounts in Step 1). The transactions would replace the transaction of the same sender and nonce sent in Step 1). Once the replacement is finished, they turn their child transactions into latent overdraft. Specifically, if a sender’s balance is $bal$ and the Ether value of transaction of nonce 2 is $v_{2}$, then $v_{1}$ is carefully crafted to enable turned latent overdraft, that is, $v_{1}<bal$ and $v_{1}+v_{2}>bal$.
@@ 5.1 Found Eviction Attacks (Selected)
In practice, many clients are found vulnerable under this exploit, including Geth $<v1.11.4$, Erigon, Besu, Nethermind, and OpenEthereum.
Exploit $XT_{6}$: Compositional indirect eviction (by multi-round valid-turned-future transactions): The latest Geth $\geq{}v1.11.4$ is patched against direct eviction attacks. To evade the protection, this eviction attack uses indirect eviction.
The attack works in two steps: 1) It first control-evicts the mempool using valid-turned-future transactions to avoid the triggering condition of its limit of transactions per sender. 2) It then conducts another run of valid-turned-future transactions, this time with transactions sent from one account, namely $1\times{}l$. It can succeed since the limit is not triggered. The overall attack cost is the fee of one transaction; compared to the single run of valid-turned-future attack, $XT_{6}$’s cost is much lower.
5.2 Found Locking Attacks (Selected)
@@
6 Performance Evaluation of mpfuzz
@@
Preliminary of GoFuzz: Upon each fuzzing iteration, GoFuzz [10] generates a bit-string of variable length, feeds it to our fuzzer code, and receives from our fuzzer a binary value of feedback, be either positive or negative. The GoFuzz further checks if this run increases the code coverage in the Geth client (as will be described, our fuzzer invokes functions in Geth’s txpool). If the feedback is positive and code coverage increases, GoFuzz adds the current bit-string to the seeds before running into the next iteration.
@@
Baseline B1: Stateless fuzzer. We implement a stateless fuzzer in the GoFuzz framework described above. 1) Given a bit string generated by GoFuzz, our fuzzer code parses it into a sequence of transactions; the length of the sequence depends on the length of the bit string. It skips a bitstring that is too short. 2) The fuzzer then sets up and initializes a Geth txpool instance with $m$ normal transactions, each sent from a distinct account, of nonce $1$, of value $1$ Ether (lower than the account balance), and of Gas price $3$ wei. 3) The fuzzer sends the transaction sequence to the initialized mempool for execution. 4) It checks the mempool state after execution: If the eviction-based test oracle (Equation 1) is met, it emits the current transaction sequence as an exploit.
@@
Baseline B2: Concrete-state-coverage fuzzer. We also build the second baseline, a mempool fuzzer that takes concrete state coverage as feedback. Specifically, in each iteration, the fuzzer appends to the current transaction sequence a new transaction. Given the $m$-slot mempool, the fuzzer tries $m$ values for senders, $m$ values for nonces, $m$ values for Gas price, and $m$ values for Ether amount. After sending to the mempool the current transaction sequence, it sorts the transactions in the mempool by senders and nonces. It then hashes the sorted transactions in the mempool and checks the presence of the hash digest in the seeds. If no hash exists, the current transaction sequence would increase the concrete-state coverage and is inserted to the corpus. The corpus is a FIFO queue where the seed inserted earlier has high priority to be de-queued.
@@
Baseline B3: Invalid-tx energy fuzzer. We build the last baseline that is similar to the concrete-state-based fuzzer except for the following: It uses as the energy the number of invalid transactions on a mempool state. The seed with more invalid transactions in the mempool has higher energy or priority to be selected for the next round of fuzzing.
@@
We implement Baseline B1 in Go/GoFuzz and implement B2 and B3 in Python. We use a test oracle that can detect DETER attacks, that is, the number of invalid transactions in a mempool state equal $m-1$.
@@
We run the three baselines and mpfuzz against a Geth client under two settings: A small setting where the mempool is resized to $6$ slots $m=6$ and all fuzzers run for two hours (if test oracle is not triggered), and a medium setting where $m=16$ and fuzzers run for $16$ hours.
@@
Table 3: Fuzzing Geth $v1.11.3$’s mempool by different approaches to detect ADAMS exploit.
Settings
B1 B2 B3 mpfuzz
6-slot-2h
timeout $54$ (min) $8$ (min)
$0.11$ (min)
16-slot-16h
timeout timeout $447$ (min)
$0.84$ (min)
@@
Table 3 reports the results under the small and medium settings. For the small setting with $6$ slots, the stateless fuzzer (baseline B1) cannot find any DETER attack in two hours, concrete-state-coverage fuzzer (baseline B2) can find a DETER exploit in 54 minutes, and the stateful fuzzer based on invalid-transaction counter finds a DETER exploit in $8$ minutes. By contrast, mpfuzz finds a DETER exploit in $0.11$ minutes.

### v3

@@ 6 Found Exploits
Table 3: ADAMS exploit patterns found by mpfuzz across Ethereum clients; $XT_{1-7}$ are eviction based, and $XT_{8-9}$ are locking based. $XT_{1-3}$ are known patterns in DETER [28], while others are new. ✗ indicates the presence of a bug, indicates the fixing of a bug after our reporting, and ✓ indicates the fixing of a bug by previous works.
$XT_{1}$ $XT_{2}$ $XT_{3}$ $XT_{4}$ $XT_{5}$ $XT_{6}$ $XT_{7}$ $XT_{8}$ $XT_{9}$
[…]
Geth $<v1.11.4$, EigenPhi
go-opera $\leq{}v1.1.3$
Erigon $\leq{}v2.42.0$
Besu $\geq{}v22.7.4$
Besu $<v22.7.4$
Nethermind $\geq{}v1.18.0$
Nethermind $<v1.18.0$
[…]
OpenEthereum $\leq{}v3.3.5$
We have run mpfuzz across a variety of Ethereum clients, including six leading execution-layer clients on the public-transaction path of Ethereum mainnet (Geth, Besu, Nethermind, Erigon, Reth, and OpenEthereum), three PBS clients (proposer-builder separation) on the mainnet’s private-transaction and bundle path (Flashbot builder $v1.11.5$ [6], EigenPhi builder [4] and bloXroute builder-ws [1]), and the clients deployed on three operational Ethereum-like networks (BSC $v1.3.8$ [2] deployed on Binance Smart Chain, go-opera $v1.1.3$ [9] on Fantom, and core-geth $v1.12.19$ [3] on Ethereum Classic).
On the six public-transaction clients, mpfuzz leads to the discovery of $22$ bugs, as listed in Table 3, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions.
Other clients, including the three PBS clients and three clients on Ethereum-like networks, are mostly forks of the Geth clients except for minor code changes [^4] For instance, go-opera adopts its own fix against $XT_{1}$ based on the fork of Geth $<v1.11.4$., and on them, mpfuzz discovered $13$ bugs of a similar nature to those found on Geth (of the historical and latest versions), as seen in Table 3.
We describe the patterns of these bugs as follows.
6.1 Found Eviction Attacks
Exploit $XT_{1}$: Direct eviction by future transactions: In this attack, given a mempool’s initial state storing normal transactions, the attacker sends future transactions at high GasPrice to evict the normal transactions. This exploit is essentially the DETER-X attack [28].
@@ 6.1 Found Eviction Attacks
In practice, Geth ($\leq{}v1.11.4$), Nethermind, Besu, and EigenPhi are vulnerable under this exploit, as in Table 3.
Exploit $XT_{2}$: Direct eviction by latent overdraft transactions: In this eviction attack, the attacker sends latent-overdraft transactions at high GasPrice to evict the normal transactions initially stored in the target mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions, denoted by $k\times{}l$. The intention is to evade the limit on the number of transactions per sender. The evasion increases the attacker cost from one pending transaction to multiple. This exploit is essentially the DETER-Z attack [28].
In practice, Geth ($\leq{}v1.11.4$), Besu, EigenPhi, and go-opera are found vulnerable under this exploit.
Exploit $XT_{3}$: Compositional direct eviction (by combining $XT_{1}$ and $XT_{2}$): In some Ethereum clients, notably Geth, the limit on the number of transactions per sender is triggered under the condition that the mempool stores enough pending transactions (e.g., more than 5120 transactions in Geth). This eviction attack combines $XT_{1}$ and $XT_{2}$ to avoid triggering the condition and evade the protection. It works by maximizing the eviction of mempool under $XT_{1}$ until it is about to trigger the condition. It then conducts $XT_{2}$ by sending latent overdraft transactions under one sender, that is, $1\times{}l$. Compared to $XT_{2}$, the combined exploit $XT_{3}$ achieves lower costs.
In practice, Geth ($\leq{}v1.11.4$) is vulnerable under this exploit where its mempool of capacity of 6144 slots triggers the limit of 16 transactions per sender when there are more than 5120 pending transactions. That is, $XT_{3}$ is configured with $l=5120$. Other clients including EigenPhi and go-opera are also vulnerable.
Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high GasPrice to evict normal transactions initially stored in the mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions. 2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high GasPrice, of high value $v_{1}$, and from the same $k$ accounts in Step 1). The transactions would replace the transaction of the same sender and nonce sent in Step 1). Once the replacement is finished, they turn their child transactions into latent overdraft. Specifically, if a sender’s balance is $bal$ and the Ether value of transaction of nonce 2 is $v_{2}$, then $v_{1}$ is carefully crafted to enable turned latent overdraft, that is, $v_{1}<bal$ and $v_{1}+v_{2}>bal$.
In practice, many clients are found vulnerable under this exploit, including Geth $<v1.11.4$, Erigon, Besu, Nethermind, OpenEthereum, EigenPhi, and go-opera.
Exploit $XT_{5}$: Indirect eviction by valid-turned-future transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high fees to evict normal transactions initially stored in the mempool. These transactions are sent by $k\times{}l$, that is, from $k$ accounts, each with $l$ transactions. For each sender, the transaction of nonce 1 has a fee, say $f_{1}$, slightly lower than the transactions of other nonces, that is, $f_{1}<f_{2}$. 2) She then sends $k$ transactions; each of the transactions is from a distinct sender from those used in Step 1) and of fee $f^{\prime}$ that $f_{1}<f^{\prime}<f_{2}$. The intent is that the $k$ transactions in Step 2) evict the transactions of nonce $1$ sent in Step 1), turning other child transactions sent in Step 1) into future transactions.
Exploit $XT_{6}$: Compositional indirect eviction (by multi-round valid-turned-future transactions): This exploit is an adaptive attack to the Geth $\geq{}v1.11.4$. Geth $\geq{}v1.11.4$ is patched against $XT_{1-2}$ and adopts the following admission policies: The mempool of $m^{\prime}=6144$ slots admits up to $py_{1}^{\prime}=1024$ future transactions. When containing more than $py_{3}^{\prime}=5120$ pending transactions, the mempool starts to limit that no more than $py_{2}^{\prime}=16$ pending transactions from the same sender can be admitted.
Exploit $XT_{6}$ works in three steps: 1) it first evicts the initial mempool of normal transactions with $m^{\prime}/py_{2}^{\prime}=384$ sequences, each of $16$ transactions sent from one unique sender, 2) it then sends $py_{1}^{\prime}/(py_{2}^{\prime}-1)=65$ transactions to evict the parent transactions and creates $py_{1}^{\prime}=1024$ future transactions in the mempool, 3) it sends one sequence of $5120$ transactions from one sender to evict the normal transactions sent in Step 1), and 4) it sends one transaction to evict the parent transaction in Step 3, leaving the mempool of just one valid transaction. Step 3) can succeed because the precondition (w.r.t. $py_{3}^{\prime}=5120$) of limiting transactions of the same sender does not hold.
In practice, Geth of all versions and the Geth forks, including Flashbot, bloXroute, BSC, and Ethereum Classic are found vulnerable under $XT_{6}$.
Exploit $XT_{7}$: Reversible evictions: Suppose a mempool of state $st$ admits an arriving transaction $tx$ by evicting an existing transaction $tx^{\prime}$, transitioning its state to $st^{\prime}$. A mempool is reversible if one sends $tx^{\prime}$ to state $st^{\prime}$, and the mempool admits $tx^{\prime}$ by evicting $tx$, transitioning its state back to $st$.
The mempool on Nethermind $<v1.18.0$ can be reversible: Minimally, suppose a two-slot mempool stores $tx_{1}$ of low GasPrice from one sender and $tx_{2}$ of medium GasPrice from another sender $A2$ and receives $tx_{3}$ as a child of $tx_{1}$ and of high GasPrice. For instance, $tx_{1}\begin{bmatrix}{A}&{1}\\ {1}&{1}\end{bmatrix}$, $tx_{2}\begin{bmatrix}{B}&{1}\\ {1}&{3}\end{bmatrix}$, $tx_{3}\begin{bmatrix}{A}&{1}\\ {2}&{5}\end{bmatrix}$. The Nethermind $<v1.18.0$ mempool would admit $tx_{3}$ and evict $tx_{2}$. After that, if one resends the evicted $tx_{2}$ back to the mempool storing $tx_{1}$ and $tx_{3}$, the Nethermind $<v1.18.0$ mempool would admit $tx_{2}$ and evict $tx_{3}$, looping back to its initial state.
An attacker observing reversible mempool mounts Exploit $XT_{7}$ to evict the mempool while bringing down attack costs. On Nethermind $<v1.18.0$, the attack works in two steps: 1) She first sends $k$ transaction sequences from $k$ senders, [^5] Nethermind does not limit transactions per sender, thus $k=1$ in actual attacks. each of $l$ transactions (i.e., $k\times{}l$). In each $l$-transaction sequence, the nonce-$1$ transaction has a low GasPrice, say $f_{1}$, while all its child transactions are of high GasPrice, evicting normal transactions initially stored in the mempool. 2) The attacker sends $k\times(l-1)$ transactions from $k\times(l-1)$ new senders, all with GasPrice slightly higher than $f_{1}$. The mempool would admit these transactions to the mempool by evicting all child transactions sent in Step 1), bringing down the attack costs.
6.2 Found Locking Attacks
@@
Exploit $XT_{9}$: Locking mempool of no turning: Because of the risk of evicting a parent transaction that could lead to turning, certain mempool is designed to restrict eviction victims to child transactions; that is, only the transaction of maximal nonce under a given sender can be evicted. For instance, given an arriving transaction $tx$, the OpenEthereum mempool finds an existing child transaction $tx^{\prime}$ with a lower GasPrice than $tx$ and evicts it to admit $tx$.
In Exploit $XT_{9}$, the attacker observes empty slots in the mempool (e.g., created by arriving blocks) and sends the same number of transactions to occupy them. The transactions are sent from one account where the child transaction has a higher GasPrice $f_{1}$ than normal transactions, while all its parent transactions have minimal fees. A normal transaction that arrives subsequently is declined because $f_{1}$ is higher than a normal transaction. Because each of its parent transactions has a low GasPrice, the mempool is locked at a low cost.
6.3 Evaluation of Found Turning Exploits
This and next subsections show the evaluation of turning attacks and locking attacks, and due to the space limit, the evaluation of other ADAMS attacks is deferred to Appendix D.
6.3.1 Evaluation in Testnet
Experimental settings: We set up our experiment platform by connecting the attacker node to the Goerli testnet. The attacker node runs an instrumented Geth client at the execution layer which sends crafted transactions to the testnet.
To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet. The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and can connect to as many neighbors as possible. The measurement node is not (directly) connected to the attacker node. When setting up our experiments, we ran the measurement node in Goerli for seven days and found the node is stabilized at $290$ neighbors.
Figure 3: Mounting $XT_{4}$ attacks on Goerli: Etherscan screenshot of the blocks generated during the attack
To do an experiment, we make the attacker node send $XT_{4}$ transactions using $384$ accounts. Each account sends $16$ valid pending transactions, followed by a replacement transaction. In total, the attacker node sends $6144$ valid transactions and $384$ replacement transactions. The GasPrice of the valid and replacement transactions are set to be $8$ Gwei and $130$ Gwei, respectively. Finally, we wrap up the experiment by waiting after the attacker node sends all messages and all replacement transactions are included in the blockchain.
Figure 3 shows the generated blocks in the experiment. We took the screenshot from etherscan.io and labeled it in red with information regarding the attack. Before the attack begins, the testnet normally utilizes $24-47\%$ of the Gas in a block. For ethics, our attacks are short (lasting four blocks).
Right after the launch of the attack, in block $xx07$, the Gas used by normal transactions drops to $3.19\%$ and the Gas used by adversarial transactions (denoted by red bars in Figure 3) is $26.88\%$. The included adversarial transactions are $384$ replacement transactions sent in the second round of $XT_{4}$. The $6144-384=5760$ child transactions sent in the first round are not included in the block. Other blocks during the attack, namely $xx08$ and $xx10$, are similar. In the third block $xx09$, $73.1\%$ Gas is spent on including $82$ normal transactions.To explain it, we inspect the raw history of transactions arriving at and logged by our measurement supernode and found that out of the $82$ included normal transactions, only $4$ are present in the log, implying the other $78$ normal transactions are private ones that were not broadcast to the measurement node. Notice that our attacks require no discovery of critical nodes as needed in [28].
6.4 Evaluation of Found Locking Exploits
6.4.1 Exploit Extension
[…]
Actual exploit $XT_{8a}$: In $XT_{8}$, the attack cost increases with the block $BasePrice$. We manually propose a method to decrease the block $BasePrice$ in Ethereum by mounting an eviction attack on the previous block. Specifically, in Ethereum (after EIP1559), given a recently produced block $i$, the block $BasePrice$ of block $i+1$ is calculated dynamically as follows:
$BasePrice(i+1)$ $=$ $BasePrice(i)*[\frac{7}{8}+\frac{1}{4}*\frac{\textit{GasUsed}(i)}{BlockLimit}]$ (10)
Therefore, if an eviction attack can persistently lower the Gas utilization of recent blocks, the current $BasePrice$ can be reduced, which decreases the costs of locking attacks on the current block. In Exploit $XT_{8a}$, we mount a series of eviction attacks first and then a locking attack against the Reth node.
Specifically, suppose in a network of a Reth node and a block validator node, the $XT_{8a}$ attacker first keeps sending eviction attacks directly to the validator node for several consecutive blocks. When observing the block $BasePrice$ drop sufficiently, the attacker mounts the regular locking attack $XT_{8}$ against the Reth node.
6.4.2 Attack Evaluation on Reth
Experimental settings: We set up an experiment platform for evaluating locking attacks, on which an attacker node is connected to a victim non-validator node, which is also connected to a workload generator node and a victim validator node. The attacker node also maintains a direct connection to the victim validator. The network topology is depicted in Figure 4. The non-validator node runs the Geth $v1.11.4$ client.
Figure 4: Experimental setup for locking attacks on Reth
Figure 5: Evaluation of locking attack $XT_{8}$ on Reth
Evaluation of $XT_{8}$: To evaluate $XT_{8}$, we set up the workload generator that sends normal transactions to the validator Geth node. The non-validator Reth node has not yet joined the network. We let the attacker node first send eviction attacks ($XT_{6}$) directly to the victim validator for $35$ consecutive blocks until it observes the block $BasePrice$ drops sufficiently to $1$ Gwei, which occurs at the $45$-th in our experiment. The Reth node then joins the network with an empty mempool. The attacker node mounts a locking attack $XT_{8}$ with transactions of price $5$ Gwei to occupy the Reth node’s mempool.
We report the total fees of normal transactions included in the blocks – the lower the fees are, the more successful the locking attack is. We also report the adversarial transaction fees as the attack cost shown in Figure 5. When the locking attack begins at the $45$-th block, the normal transaction fees increase because of the empty mempool on the Reth node which accepts some normal transactions in addition to the adversarial locking transactions. Then, after three blocks, the normal transaction fees quickly drop to near-zero Ether, showing the success of locking attacks. Interestingly, unlike the eviction attacks that cause low Gas utilization per block, the locking attacks just reduce the total Ether fees without reducing Gas utilization (i.e., the blocks produced under locking attacks from height $45$ to $60$ use the Gas of almost $100\%$ block limit). Due to the high Gas utilization, block $BasePrice$ keeps increasing during the locking attack. When reaching the $60$-th block, the $BasePrice$ is higher than the maximal transaction price tolerable by the attacker. In our experiment, the locking attacker stops at $60$-th block, and the normal transaction fees immediately recover to the “normal” level. In practice, the attacker can repeat sending eviction attacks to reduce the $BasePrice$ before mounting the locking attack again.
7 Evaluation of mpfuzz
7.1 Performance
@@
Preliminary of GoFuzz: Upon each fuzzing iteration, GoFuzz [16] generates a bit-string of variable length, feeds it to our fuzzer code, and receives from our fuzzer a binary value of feedback, be either positive or negative. The GoFuzz further checks if this run increases the code coverage. If the feedback is positive and code coverage increases, GoFuzz adds the current bit-string to the seeds before running into the next iteration.
@@
Baseline B1: Stateless fuzzer. We implement a stateless fuzzer in the GoFuzz framework described above. 1) Given a bit string generated by GoFuzz, our fuzzer code parses it into a sequence of transactions; the length of the sequence depends on the length of the bit string. It skips a bitstring that is too short. 2) The fuzzer sets up and initializes a Geth txpool instance with $m$ normal transactions, each sent from a distinct account, of nonce $1$, of value $1$ Ether (lower than the account balance), and of GasPrice $3$ wei. 3) The fuzzer sends the transaction sequence to the initialized mempool for execution. 4) It checks the mempool state after execution: If the test oracle is met, it emits the current transaction sequence as an exploit.
@@
Baseline B2: Concrete-state-coverage fuzzer. We build the second baseline, a mempool fuzzer that takes concrete state coverage as feedback. Specifically, in each iteration, the fuzzer appends a new transaction to the current transaction sequence. Given the $m$-slot mempool, the fuzzer tries $m$ values for senders, nonces, GasPrice, and Ether amount. After sending the current transaction sequence to the mempool, it sorts the transactions in the mempool by senders and nonces. It then hashes the sorted transactions in the mempool and checks the presence of the hash digest in the seeds. If no hash exists, the current transaction sequence would increase the concrete-state coverage and is inserted to the corpus. The corpus is a FIFO queue where the seed inserted earlier has high priority to be de-queued.
@@
Baseline B3: Invalid-tx energy fuzzer. Baseline B3 is similar to B2 except that it uses as the energy the number of invalid transactions on a mempool state. The seed with more invalid transactions in the mempool has higher energy or priority to be selected for the next round of fuzzing.
Baseline B4: feedback of no promising states: We design Baseline B4 for an ablation study on the effectiveness of using promising states as feedback in improving fuzzing performance. B4 is the same with mpfuzz except that it removes the state promising-ness from its feedback. Specifically, the feedback formula in B4 includes only state coverage ($st\_coverage$) and excludes state promisingness ($st\_promising$ as in Equation 7).
@@
We implement Baseline B1 in Go/GoFuzz and implement B2, B3 and B4 in Python. We use a test oracle that can detect DETER attacks, that is, the number of invalid transactions in a mempool state equal $m-1$.
@@
Experimental settings of fuzzing: We run the four baselines and mpfuzz against a MUT in a Geth client reconfigured under two settings: A small setting where the mempool is resized to $6$ slots $m=6$ and all fuzzers run for two hours (if test oracle is not triggered), and a medium setting where $m=16$ and fuzzers run for $16$ hours. The fuzzing experiments are run on a local machine with an Intel i7-7700k CPU of $4$ cores and $64$ GB RAM.
@@
Table 4: Fuzzing Geth $v1.11.3$’s mempool (in minutes) by different approaches to detect Exploit $XT_{3}$.
Settings B1 B2 B3 B4 mpfuzz
$6$slot-$2$h Timeout $54$ $8$ $1.22$ $0.03$
$16$slot-$16$h Timeout Timeout $447$ Timeout $0.06$
@@
Results: Table 4 reports the results under the small and medium MUT settings. For the small MUT, B1 cannot find any DETER attack in two hours. B2, B3 and B4 can find $XT_{3}$ in 54 minutes, $8$ minutes and $1.22$ minutes respectively. By contrast, mpfuzz finds the same exploit in $0.03$ minutes.